# SAKE: Towards Editing Auditory Attribute Knowledge of Large Audio-Language Models

### v1

@@ Appendix F Detailed Results under Sequential Editing
FT (LLM) 0 100.00 100.00 44.56 64.00 18.00
1 82.00 74.67 41.97 70.00 16.00
2 72.00 61.33 41.97 64.00 18.00
3 70.00 55.33 42.49 64.00 14.00
4 66.00 51.33 45.08 62.00 20.00
5 56.00 43.33 40.41 66.00 16.00
FT (Audio) 0 100.00 82.67 74.09 100.00 42.00
1 100.00 79.33 72.54 100.00 40.00
2 100.00 77.33 71.50 100.00 40.00
3 100.00 76.00 70.98 100.00 36.00
4 100.00 76.00 70.98 100.00 34.00
5 100.00 75.33 66.84 100.00 40.00
MEND 0 88.00 88.67 64.25 86.00 20.00
1 76.00 72.67 68.91 82.00 20.00
2 66.00 68.67 65.28 66.00 16.00
3 68.00 68.00 64.25 72.00 22.00
4 64.00 66.67 58.55 70.00 14.00
5 64.00 55.33 57.51 70.00 18.00
KE 0 72.00 70.67 41.97 54.00 12.00
1 34.00 27.33 29.02 40.00 8.00
2 22.00 17.33 22.28 32.00 2.00
3 10.00 17.33 18.65 24.00 4.00
4 8.00 10.67 13.47 14.00 2.00
5 14.00 9.33 13.47 12.00 4.00
UnKE 0 100.00 98.00 47.67 66.00 26.00
1 54.00 54.00 37.82 62.00 10.00
2 50.00 50.00 38.86 54.00 24.00
3 42.00 34.67 40.93 50.00 22.00
4 38.00 29.33 29.53 52.00 24.00
5 40.00 26.00 33.16 36.00 16.00
IE-IKE 0 10.00 6.00 77.20 38.00 26.00
1 14.00 7.33 72.02 40.00 22.00
2 4.00 6.67 55.44 34.00 26.00
3 14.00 6.00 42.49 34.00 20.00
4 8.00 6.67 27.46 20.00 10.00
5 4.00 4.00 20.73 14.00 8.00
I-IKE 0 10.00 7.33 76.17 48.00 26.00
1 10.00 7.33 77.20 44.00 22.00
2 10.00 8.00 75.13 48.00 26.00
3 10.00 8.00 75.65 52.00 20.00
4 10.00 8.67 75.65 42.00 26.00
5 10.00 6.67 77.72 44.00 22.00
Figure 4: An example of a successful editing result by FT (Audio) on Qwen2-Audio.
Figure 5: An example of a degenerated editing result by MEND on DeSTA2.5-Audio.
Appendix G Case Study
@@
Figure 4 shows the result of a successful editing example by FT (Audio) on Qwen2-Audio, where we edit the model’s perception of speaker emotion from “fearful” to “sad.” After editing, we observe that the model’s outputs for both reliability and generality are successfully updated to “sad.” At the same time, both audio locality and text locality remain intact, which shows that the original knowledge of the model is preserved. For portability, we observe that after the edit, the model can also perform reasoning, changing the prediction from “preparing for a high-stakes exam with anxiety” to “a person crying after a breakup.” This demonstrates that the interconnected knowledge is also updated during reasoning, and the edited model can apply this knowledge in new contexts.
Figure 5 shows a degenerated example from sequential editing on DeSTA2.5-Audio with MEND. After applying multiple edits in a row, we can see that the model collapses and produces incoherent outputs such as repeated characters and newline symbols. This illustrates how sequential edits can accumulate interference and destabilize the internal representations of the model. Unlike the single-edit scenario, where the change is targeted and localized, multiple edits interact in unpredictable ways, leading to corruption of reliability, loss of generality, and failure of both locality and portability. To be practical in real-world scenarios, however, an editing method must be capable of supporting many edits simultaneously while ensuring that unrelated edits do not interfere with one another. This failure case thus underscores a key limitation of current approaches.

### v2

@@ Appendix G Detailed Results under Sequential Editing
FT (LLM) 0 100.00 100.00 51.30 68.00 18.00
1 82.00 74.67 46.11 66.00 16.00
2 72.00 60.67 44.56 64.00 18.00
3 70.00 54.00 45.60 66.00 14.00
4 66.00 50.67 47.15 66.00 18.00
5 56.00 43.33 40.93 58.00 16.00
FT (Audio) 0 100.00 82.67 82.90 100.00 40.00
1 100.00 78.67 80.83 100.00 40.00
2 100.00 77.33 79.27 100.00 38.00
3 100.00 75.33 78.76 100.00 34.00
4 100.00 76.00 77.72 100.00 34.00
5 98.00 75.33 74.09 100.00 40.00
MEND 0 88.00 88.67 72.02 86.00 18.00
1 76.00 72.67 73.58 82.00 20.00
2 66.00 68.67 69.43 66.00 18.00
3 68.00 68.00 69.43 72.00 18.00
4 62.00 66.67 63.73 68.00 12.00
5 60.00 54.67 62.18 68.00 18.00
KE 0 72.00 70.67 43.01 50.00 14.00
1 28.00 27.33 30.05 38.00 4.00
2 16.00 16.67 24.87 34.00 2.00
3 10.00 14.00 19.17 22.00 0.00
4 8.00 8.67 15.03 18.00 2.00
5 6.00 7.33 13.99 12.00 2.00
UnKE 0 96.00 96.67 47.15 56.00 20.00
1 52.00 50.67 38.34 58.00 6.00
2 42.00 36.00 37.82 58.00 26.00
3 34.00 28.00 36.78 52.00 12.00
4 28.00 27.33 30.57 50.00 18.00
5 30.00 20.00 27.46 50.00 22.00
IE-IKE 0 10.00 6.00 83.42 40.00 26.00
1 14.00 7.33 78.76 42.00 22.00
2 4.00 6.67 61.14 34.00 26.00
3 14.00 5.33 47.15 26.00 18.00
4 6.00 6.00 30.57 18.00 10.00
5 4.00 3.33 21.24 12.00 8.00
I-IKE 0 10.00 7.33 88.60 54.00 24.00
1 10.00 7.33 90.16 48.00 22.00
2 10.00 8.00 88.08 48.00 22.00
3 10.00 8.00 88.08 48.00 20.00
4 10.00 8.00 88.08 46.00 22.00
5 10.00 6.67 89.12 46.00 22.00
WISE 0 100.00 100.00 58.55 84.00 14.00
1 92.00 90.00 58.03 82.00 10.00
2 92.00 78.67 62.18 76.00 12.00
3 98.00 82.67 61.14 70.00 14.00
4 94.00 80.67 61.65 76.00 18.00
5 96.00 80.00 63.73 66.00 18.00
Table 13: Original result of the four metrics of different editing methods on AF under sequential editing. For generality and audio locality, we present the averaged results. (%)
FT (LLM) 0 100.00 94.67 82.90 90.00 24.00
1 82.00 76.67 79.79 90.00 24.00
2 78.00 70.67 79.27 90.00 32.00
3 72.00 65.33 79.79 94.00 28.00
4 70.00 63.33 78.24 92.00 26.00
5 60.00 58.67 77.20 88.00 32.00
FT (Audio) 0 100.00 70.00 87.56 100.00 44.00
1 100.00 68.00 84.97 100.00 46.00
2 100.00 68.00 80.31 100.00 46.00
3 98.00 64.00 76.68 100.00 44.00
4 98.00 62.00 76.17 100.00 44.00
5 94.00 60.67 75.65 100.00 40.00
MEND 0 100.00 94.67 86.53 90.00 26.00
1 98.00 92.00 82.38 96.00 26.00
2 98.00 90.67 83.94 92.00 30.00
3 98.00 92.67 81.35 96.00 28.00
4 98.00 92.67 80.83 94.00 26.00
5 94.00 85.33 80.83 94.00 26.00
KE 0 68.00 72.67 52.33 66.00 24.00
1 60.00 54.67 39.38 48.00 20.00
2 38.00 30.00 24.87 44.00 18.00
3 24.00 26.00 22.80 28.00 12.00
4 16.00 16.00 18.65 22.00 8.00
5 10.00 13.33 12.95 16.00 2.00
UnKE 0 98.00 88.67 68.39 44.00 22.00
1 88.00 77.33 61.65 50.00 24.00
2 80.00 69.33 59.58 32.65 18.00
3 82.00 66.67 56.48 12.24 12.00
4 80.00 64.00 53.36 16.00 14.00
5 72.00 58.00 48.70 8.00 16.00
IE-IKE 0 56.00 61.33 37.82 84.00 32.00
1 52.00 52.00 39.38 82.00 28.00
2 50.00 47.33 40.93 76.00 30.00
3 42.00 40.67 39.38 78.00 34.00
4 38.00 40.00 38.34 82.00 30.00
5 38.00 41.33 42.49 74.00 30.00
I-IKE 0 22.00 9.33 84.97 74.00 38.00
1 22.00 7.33 86.53 74.00 34.00
2 24.00 7.33 87.05 80.00 34.00
3 22.00 5.33 87.05 80.00 36.00
4 22.00 6.00 87.05 76.00 32.00
5 18.00 6.67 86.01 78.00 32.00
WISE 0 100.00 74.67 88.08 90.00 24.00
1 84.00 42.00 89.12 94.00 22.00
2 64.00 32.67 89.12 92.00 22.00
3 42.00 22.00 89.12 90.00 22.00
4 34.00 18.67 87.56 90.00 22.00
5 26.00 12.67 85.49 90.00 22.00
Appendix H More Discussions of Intra-attribute Knowledge Entanglement
In Sec. 5.2, we observe that Audio Locality Type 2 is the most difficult to preserve. We hypothesize that this failure is related to the fact that intra-attribute distinctions are often more tightly coupled than inter-attribute ones, making them more vulnerable to interference during editing. In particular, both acoustic and semantic factors may jointly contribute to this phenomenon.
Acoustic Entanglement
Compared to inter-attribute distinctions (e.g., emotions versus animal sounds), labels within the same attribute often share overlapping acoustic characteristics. For example, within the emotion attribute, “sad” and “fearful” speech may both exhibit lower energy, increased tension, or slower speaking rates. Such shared prosodic and spectral patterns suggest that modifying knowledge associated with one label may unintentionally affect closely related labels, making intra-attribute knowledge harder to preserve during editing.
Semantic Entanglement
As shown in Figure 3, FT (LLM) performs substantially worse than FT (Audio) on Audio Locality Type 2. This gap suggests that intra-attribute distinctions are more susceptible to interference when updates are applied to the LLM backbone. A plausible explanation is that these distinctions may also rely on fine-grained semantic cues that are closely distributed within the models. Directly modifying LLM parameters may therefore increase the risk of conflating related labels.
In summary, current editing methods do not explicitly model the fine-grained structure of intra-attribute knowledge. Our results suggest that both acoustic overlap and semantic coupling may contribute to the observed failures in Audio Locality Type 2. Notably, identifying intra-attribute knowledge entanglement as a key challenge is itself an important insight enabled by our comprehensive benchmark design. Developing editing approaches that can better isolate and preserve intra-attribute distinctions remains an important direction for future work.
Appendix I Detailed Breakdown of Portability Evaluation
Given that the scope of potential related knowledge is vast and cannot be entirely enumerated, SAKE addresses this by aggregating portability evaluations across a diverse set of editing instances. Specifically, within our dataset, multiple editing instances may share the same editing pair $(y_{o},y_{e})$; however, they are assessed using different portability questions that target a wide variety of related concepts of the edited attribute. To provide a granular analysis of knowledge portability across these concepts, we detail the specific categories covered for each auditory attribute and present the performance breakdown under single editing.
I.1 Taxonomy of Knowledge Categories
We categorize portability questions according to the underlying reasoning concepts they involve. The resulting distribution of knowledge categories is illustrated in Figure 5.
Animal Sounds
  • Behavior: Typical behavior of the animal (e.g., purring, chewing cud).
  • Care Item: Required items for caring for the animal (e.g., litter box, leash).
  • Diet: Dietary classification of the animal.
  • Family: The closest animal in terms of biological taxonomy.
  • Locomotion: The way the animal moves.
  • Physical: Physical traits of the animal (e.g., claws, hooves).
  • Reproduction: The reproductive method of the animal.
  • Vocalization: The characteristic vocalization of the animal.
Figure 5: Distribution of portability knowledge categories associated with edited auditory attributes.
Speaker Emotion
  • Descriptive Sentence: A sentence reflecting the speaker’s emotional state.
  • Facial Expression: The facial expression representing the emotion.
  • Scenario: The scenario or situation that matches the speaker’s emotion.
  • Social Interaction: The appropriate social interaction as a response based on the emotion.
Speaker Gender
  • History: A historical figure who shares the same gender as the speaker.
  • Celebrity: A celebrity who shares the same gender as the speaker.
  • Cloth: Traditional clothing associated with the speaker’s gender in the relevant cultural context.
  • Title: Formal title of the speaker (e.g., Mr., Ms.).
  • Vocal: Vocal ranges that align with the speaker’s gender.
Spoken Language
  • ISO: The ISO language code of the spoken language.
  • City: A city in a country where the spoken language is recognized as official.
  • Dish: A dish originating from a country where the language is official.
  • History: A historical figure from a country where the language is official.
  • Language Family: The language family to which the language belongs.
  • Literary: Books or literary works originally authored in the language.
  • Official Language: The country that recognizes the spoken language as official.
  • Test: A test used to evaluate proficiency in the language.
  • Translation: The translation of an English word into the spoken language.
I.2 Performance Breakdown
Tables 14, 15, 16, and 17 detail the portability breakdown under single editting for Animal Sounds, Speaker Emotion, Speaker Gender, and Spoken Language attributes, respectively. The analysis reveals a mixed state of reasoning propagation. Rather than a binary outcome where all related concepts succeed or fail simultaneously, performance varies by category. For instance, within the Animal Sound attribute, FT (Audio) shows a disparity in knowledge propagation, achieving 45.95% accuracy on “reproduction” compared to only 7.89% on “family.” Overall, instances where related concepts fail to update still constitute the majority, indicating that current editing methods struggle to consistently propagate knowledge across different reasoning dimensions.
We further emphasize that the LALMs used in this study already possess the world knowledge required by our portability dataset (Yang et al., 2025b), whose connected attribute knowledge is drawn from the SAKURA benchmark. Therefore, observed portability failures can be attributed to ineffective edit propagation to relevant knowledge, rather than a lack of underlying world knowledge.
Table 14: Breakdown of portability performance across different knowledge categories for the Animal Sound attribute. (%)
Model Method Behavior Care Item Diet Family Locomotion Physical Reproduction Vocalization
DeSTA FT (LLM) 2.70 23.68 21.05 2.63 68.57 13.51 16.22 15.00
FT (Audio) 24.32 39.47 21.05 5.26 25.71 32.43 45.95 40.00
KE 10.81 39.47 5.26 26.32 34.29 24.32 8.11 37.50
MEND 5.41 36.84 39.47 10.53 34.29 10.81 18.92 7.50
UnKE 21.62 55.26 26.32 5.26 28.57 16.22 18.92 45.00
I-IKE 83.78 92.11 68.42 47.37 85.71 78.38 78.38 92.50
IE-IKE 32.43 57.89 42.11 13.16 51.43 37.84 35.14 80.00
WISE 2.70 15.79 15.79 0.00 17.14 10.81 27.03 2.50
Qwen FT (LLM) 21.62 26.32 13.16 23.68 31.43 21.62 10.81 30.00
FT (Audio) 56.76 71.05 36.84 28.95 37.14 51.35 43.24 60.00
KE 24.32 31.58 13.16 23.68 31.43 10.81 8.11 40.00
MEND 18.92 42.11 15.79 15.79 11.43 16.22 18.92 30.00
UnKE 40.54 21.05 18.42 10.53 0.00 8.11 10.81 50.00
I-IKE 29.73 34.21 21.05 18.42 31.43 10.81 5.41 40.00
IE-IKE 27.03 50.00 21.05 21.05 22.86 18.92 5.41 42.50
WISE 18.92 18.42 0.00 0.00 2.86 0.00 0.00 12.50
AF FT (LLM) 35.14 39.47 26.32 18.42 20.00 32.43 32.43 27.50
FT (Audio) 54.05 60.53 28.95 28.95 40.00 48.65 51.35 57.50
KE 37.84 39.47 23.68 23.68 17.14 32.43 24.32 27.50
MEND 35.14 47.37 21.05 28.95 22.86 32.43 29.73 27.50
UnKE 35.14 26.32 23.68 23.68 17.14 21.62 35.14 37.50
I-IKE 40.54 52.63 23.68 26.32 31.43 35.14 43.24 47.50
IE-IKE 48.65 47.37 28.95 31.58 34.29 35.14 45.95 55.00
WISE 29.73 47.37 21.05 26.32 25.71 29.73 35.14 25.00
Table 15: Breakdown of portability performance across different knowledge categories for the Speaker Emotion attribute. (%)
Model Method Descriptive Sentence Facial Expression Scenario Social Interaction
DeSTA FT (LLM) 24.32 19.48 17.11 27.40
FT (Audio) 44.59 46.75 47.37 45.21
KE 17.57 14.29 11.84 20.55
MEND 22.97 23.38 21.05 27.40
UnKE 16.22 24.68 13.16 19.18
I-IKE 72.97 51.95 68.42 68.49
IE-IKE 45.95 23.38 19.74 26.03
WISE 1.35 1.30 5.26 2.78
Qwen FT (LLM) 18.92 16.88 10.53 30.14
FT (Audio) 47.30 44.16 35.53 42.47
KE 21.62 15.58 22.37 31.51
MEND 24.32 15.58 22.37 31.51
UnKE 13.51 12.99 10.53 28.77
I-IKE 22.97 16.88 15.79 32.88
IE-IKE 21.62 18.18 18.42 28.77
WISE 22.97 10.39 19.74 34.25
AF FT (LLM) 14.86 18.18 21.05 17.81
FT (Audio) 36.49 42.86 46.05 27.40
KE 14.86 20.78 22.37 19.18
MEND 17.57 18.18 19.74 17.81
UnKE 12.16 16.88 21.05 20.55
I-IKE 16.22 25.97 27.63 20.55
IE-IKE 24.32 23.38 25.00 30.14
WISE 12.16 20.78 22.37 17.81
Table 16: Breakdown of portability performance across different knowledge categories for the Speaker Gender attribute. (%)
Model Method History Celebrity Cloth Title Vocal
DeSTA FT (LLM) 6.35 2.99 3.33 6.56 12.24
FT (Audio) 85.71 80.60 71.67 93.44 81.63
KE 7.94 4.48 10.00 4.92 12.24
MEND 4.76 2.99 3.33 3.28 10.20
UnKE 4.76 4.48 8.33 4.92 8.16
I-IKE 80.95 53.73 70.00 78.69 83.67
IE-IKE 36.51 23.88 18.33 45.90 40.82
WISE 33.33 22.39 31.67 36.07 24.49
Qwen FT (LLM) 47.62 11.94 18.33 32.79 18.37
FT (Audio) 90.48 35.82 40.00 88.52 57.14
KE 61.90 28.36 36.67 40.98 20.41
MEND 63.49 16.42 25.00 52.46 20.41
UnKE 57.14 32.84 36.67 34.43 26.53
I-IKE 77.78 28.36 31.67 55.74 28.57
IE-IKE 65.08 28.36 33.33 36.07 36.73
WISE 68.25 22.39 30.00 59.02 22.45
AF FT (LLM) 76.19 40.30 51.67 49.18 38.78
FT (Audio) 82.54 43.28 55.00 81.97 53.06
KE 76.19 41.79 48.33 49.18 40.82
MEND 74.60 41.79 48.33 50.82 40.82
UnKE 77.78 32.84 51.67 60.66 42.86
I-IKE 73.02 44.78 60.00 65.57 46.94
IE-IKE 77.78 44.78 60.00 67.21 57.14
WISE 71.43 38.81 50.00 44.26 38.78
Table 17: Breakdown of portability performance across different knowledge categories for the Spoken Language attribute. (%)
Model Method Iso City Dish History Language Family Literary Official Language Test Translation
DeSTA FT (LLM) 10.71 41.18 37.14 38.89 14.81 2.86 23.53 30.77 40.62
FT (Audio) 67.86 58.82 74.29 66.67 74.07 54.29 58.82 41.03 81.25
KE 14.29 23.53 22.86 30.56 14.81 11.43 26.47 15.38 40.62
MEND 14.29 23.53 28.57 22.22 11.11 14.29 5.88 41.03 46.88
UnKE 32.14 17.65 5.71 16.67 18.52 8.57 5.88 23.08 28.12
I-IKE 82.14 70.59 77.14 72.22 59.26 48.57 61.76 66.67 84.38
IE-IKE 57.14 26.47 37.14 22.22 11.11 25.71 23.53 23.08 65.62
WISE 78.57 20.59 17.14 2.78 37.04 2.86 41.18 12.82 34.38
Qwen FT (LLM) 3.57 55.88 34.29 36.11 3.70 5.71 26.47 33.33 50.00
FT (Audio) 75.00 44.12 40.00 38.89 44.44 31.43 64.71 28.21 62.50
KE 3.57 26.47 28.57 19.44 14.81 17.14 14.71 30.77 46.88
MEND 17.86 29.41 25.71 13.89 18.52 17.14 26.47 23.08 46.88
UnKE 64.29 35.29 14.29 22.22 51.85 20.00 32.35 15.38 43.75
I-IKE 10.71 35.29 20.00 19.44 18.52 14.29 17.65 23.08 46.88
IE-IKE 7.14 26.47 25.71 13.89 29.63 14.29 8.82 25.64 50.00
WISE 78.57 44.12 45.71 36.11 40.74 5.71 44.12 33.33 59.38
AF FT (LLM) 17.86 20.59 14.29 25.00 29.63 34.29 32.35 33.33 37.50
FT (Audio) 92.86 32.35 51.43 55.56 55.56 31.43 41.18 35.90 46.88
KE 28.57 20.59 25.71 25.00 37.04 31.43 17.65 30.77 37.50
MEND 21.43 20.59 17.14 27.78 25.93 34.29 20.59 33.33 40.62
UnKE 10.71 20.59 17.14 22.22 18.52 31.43 23.53 28.21 40.62
I-IKE 25.00 26.47 25.71 25.00 29.63 37.14 23.53 30.77 37.50
IE-IKE 57.14 32.35 25.71 33.33 33.33 34.29 35.29 25.64 40.62
WISE 42.86 23.53 22.86 25.00 25.93 25.71 26.47 25.64 37.50
Appendix J Case Study
@@ I.2 Performance Breakdown
Example of successful editing.
Figure 6 shows the result of a successful editing example by FT (Audio) on Qwen, where we edit the model’s perception of speaker emotion from “fearful” to “sad.” After editing, we observe that the model’s outputs for both reliability and generality are successfully updated to “sad.” At the same time, both audio locality and text locality remain intact, which shows that the original knowledge of the model is preserved. For portability, we observe that after the edit, the model can also perform reasoning, changing the prediction from “preparing for a high-stakes exam with anxiety” to “a person crying after a breakup.” This demonstrates that the interconnected knowledge is also updated during reasoning, and the edited model can apply this knowledge in new contexts.
Degeneration Analysis.
Figure 7 shows a degenerated example from sequential editing on DeSTA with MEND. After applying multiple edits in a row, we can see that the model collapses and produces incoherent outputs such as repeated characters and newline symbols. For example, some of the output will become ”happy catholic catholic catholic catholic……” or ”dogdogdogdogdog….” This illustrates how sequential edits can accumulate interference and destabilize the internal representations of the model. Unlike the single-edit scenario, where the change is targeted and localized, multiple edits interact in unpredictable ways, leading to corruption of reliability, loss of generality, and failure of both locality and portability. To be practical in real-world scenarios, however, an editing method must be capable of supporting many edits simultaneously while ensuring that unrelated edits do not interfere with one another. This failure case thus underscores a key limitation of current approaches.
Quantitatively, we detect repetitive degeneration patterns in model outputs using regular expressions for each editing method and LALM under sequential editing. The frequency of degeneration is summarized in Table 18. These statistics are consistent with the trends in Figure 4, where MEND exhibits severe degeneration on DeSTA, KE on all models, and UnKE on AF.
Table 18: Ratio (%) of degenerated outputs of each editing methods on each LALM under sequential editing.
Model FT (LLM) FT (Audio) KE MEND UnKE WISE IE-IKE I-IKE
DeSTA 0.30 1.35 47.77 44.25 2.23 0.14 0.00 0.78
Qwen 0.34 0.20 11.97 0.85 2.30 0.20 1.28 0.34
AF 0.00 0.00 9.67 0.00 6.02 0.00 0.00 0.00
Figure 6: An example of a successful editing result by FT (Audio) on Qwen.
Figure 7: An example of a degenerated editing result by MEND on DeSTA.